The volume of a cone varies jointly as the square of its radius and its height. If the volume of a cone is 32pi cubic inches when the radius is 4 inches and the height is 6 inches, find the volume of a cone when the radius is 3 inches and the height is 5 inches.
Answer by mukhopadhyay(490)

V=kr^2h
32pi = k(16)(6)
=> k = pi/3
Volume V = pi(r^2h)/3;
V = pi(45)/3 (using specs from the question)
=> V = 15pi
Answer: 15pi Cubic Inches